OpenStudy (anonymous):

Add. 5 and 1/8 + 5\6 + 5\12

OpenStudy (anonymous):

5+(1/8) = (41/8) (41/8)+(5/6)+(5/12) = (51/8)

OpenStudy (anonymous):

thats not one of the awnsers

OpenStudy (anonymous):

read my message to you :)

OpenStudy (anonymous):

yea i saw idk how to but it on top and bottom "6/6"